# Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

### v2

@@ Abstract
The advancement of general-purpose artificial intelligence relies on large language models (LLMs) that excel across a wide range of tasks, from structured reasoning to creative generation. However, post-training methods like Supervised Fine-Tuning (SFT) often struggle with generalization, favoring memorization over transferable learning. In this work, we introduce Omni-Thinker, a unified reinforcement learning (RL) framework that enhances LLM performance across diverse tasks by combining rule-based verifiable rewards with generative preference signals via LLM-as-a-Judge evaluations. Our approach enables consistent optimization across task types and scales RL-based training to subjective domains. We further investigate training strategies, demonstrating that a curriculum-based progression that orders tasks from structured to open-ended improves performance and reduces forgetting. Experimental results across four domains reveal that curriculum learning improves performance by 5.2% over joint training and 9.1% over model merging. These results highlight the importance of task-aware sampling and hybrid supervision in scaling RL-based post-training for general-purpose LLMs.
@@ 1 Introduction
Figure 1: OMNI-THINKER Framework Overview. A unified multi-task RL framework for training a single LLM across structured and open-ended tasks. (1) Backward Transfer quantifies task forgettability and informs (2) a curriculum that schedules tasks from least to most forgettable. (3) Sampled tasks construct prompts. (4) The LLM generates completions, which are (5) scored by hybrid rewards: verifiable (e.g., test case execution), generative (e.g., LLM-as-a-Judge), and auxiliary (e.g., format adherence), among others. (6) The policy is updated via Multi-Task GRPO. The process repeats via iterative updates (bottom loop).
As Large Language Models (LLMs) (Hurst et al., 2024; Liu et al., 2024; Dubey et al., 2024; Yang et al., 2024) evolve into general-purpose agents with applications ranging from creative writing (Marco et al., 2025) to robotics (Gemini-Robotics-Team et al., 2025), the challenge of effective post-training becomes increasingly critical. While Supervised Fine-Tuning (SFT) is the dominant paradigm for adapting LLMs to downstream tasks, it often encourages memorization rather than robust generalization, especially on tasks that differ from the training distribution or require flexible, context-dependent reasoning (Chu et al., 2025).
@@
Reinforcement Learning (RL) has emerged as a promising path to improve generalization, especially in structured tasks like math and coding, where verifiable, rule-based rewards can be effective (DeepSeek-AI et al., 2025; Luo et al., 2025; Kimi-Team et al., 2025). Methods such as Group Relative Policy Optimization (GRPO) have shown that even coarse signals can guide LLMs toward structured, chain-of-thought responses (Shao et al., 2024). However, most RL methods are tailored to deterministic, easily verifiable tasks, limiting their utility in open-ended settings such as question answering and creative writing. Moreover, training LLMs across multiple tasks remains challenging because it requires optimizing for diverse forms of feedback signals, including binary correctness checks in structured tasks and subjective, preference-based judgments in generative ones.
@@
To address this, we propose Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision across tasks. Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks. We further show that curriculum training, which sequences tasks from structured to open-ended based on Backward Transfer (BWT) measurements of task forgettability, significantly improves generalization and reduces forgetting. Across four domains, it achieves average gains of 5.2% and 9.1% over joint training and model merging, respectively.
Our key contributions are: (1) We propose Omni-Thinker, a unified framework that supports both verifiable and generative supervision under a single policy, scaling RL across four diverse domains. (2) We show that LLM-as-a-Judge enables scalable reward signals for open-ended tasks, extending GRPO beyond rule-based domains. (3) We demonstrate that not all multi-task training strategies are equally effective. Curriculum-based task scheduling outperforms uniform sampling and improves generalization in multi-task RL.
2 Related Work
Large Language Models and Multi-Task Learning
Large Language Models and Reinforcement Learning
3 Problem Formulation
We consider the problem of training a single language model policy to perform well across a diverse set of tasks, including mathematical reasoning, code generation, question answering, and creative writing. Let $\mathcal{T}=\{T_{1},\ldots,T_{K}\}$ be a collection of $K$ tasks, where each task $T_{k}$ has a dataset $\mathcal{D}_{k}$ of input-output pairs $(q_{k},o^{*}_{k})$, where $q_{k}\in\mathcal{Q}_{k}$ is an input prompt, and $o^{*}_{k}\in\mathcal{O}^{*}_{k}$ is the corresponding ground-truth output response.
The model, parameterized by $\theta$, defines a conditional distribution $\pi_{\theta}(y_{k}\mid x_{k})$. Supervised fine-tuning (SFT) aims to maximize the likelihood of ground-truth outputs by minimizing:
$\min_{\theta}\sum_{k=1}^{K}\mathbb{E}_{(x_{k},y_{k})\sim\mathcal{D}_{k}}\left[\sum_{t=1}^{T}-\log\pi_{\theta}(o^{*}_{k,t}\mid q_{k},o^{*}_{k,<t})\right],$ (1)
where $o^{*}_{k,t}$ denotes the $t$-th token of $o^{*}_{k}$, and $T$ is the number of tokens in the output.
While effective in-domain, SFT often overfits to training data and fails to generalize, particularly when task formats vary. Recent studies (Chu et al., 2025) show that reinforcement learning (RL) with outcome-based rewards better promotes cross-task generalization. Motivated by this, we adopt a multi-task RL (MTRL) formulation with task-specific reward functions $R_{k}$. The goal is to learn a unified policy $\pi_{\theta}$ that maximizes the expected reward over the task distribution:
@@ 3 Problem Formulation
$\max_{\theta}\;\mathcal{J}(\theta)=\mathbb{E}_{T_{k}\sim P(\mathcal{T})}\left[\mathbb{E}_{q_{k}\sim\mathcal{D}_{k},o_{k}\sim\pi_{\theta}(\cdot\mid x_{k})}\left[R_{k}(o_{k})\right]\right],$ (2)
where $P(\mathcal{T})$ is the task sampling distribution, which determines task exposure during training, and $R_{k}$ is the task-specific reward function. This objective trains a single policy, $\pi_{\theta}$, that balances performance across tasks, enabling knowledge transfer, mitigating negative transfer, and improving generalization across domains.
4 Methodology
4.1 Omni-Thinker Framework Overview
We introduce Omni-Thinker, a unified post-training framework that aligns LLMs to a wide range of tasks via reinforcement learning, extending RLVR beyond structured domains like math and code to include general and open-ended tasks.
@@ 4.1 Omni-Thinker Framework Overview
Together, these strategies enable consistent RL training across structured and generative tasks within a unified optimization framework.
4.2 Multi-Task Group Relative Policy Optimization
We extend the GRPO (Guo et al., 2025) algorithm to the multi-task setting by jointly optimizing over task-specific reward signals and reference policies. For each input prompt $q_{k}$, GRPO samples a group of outputs $\{o_{k,1},o_{k,2},\cdots,o_{k,G}\}$ from the old policy $\pi_{\theta_{old}}$. At each generation step $t$, the partial output is denoted $o_{k,i,<t}$, where $o_{k,i,<t}=[o_{k,i,1},\dots,o_{k,i,t-1}]$, that is, the input concatenated with the sequence of generated tokens up to step $t-1$. A reward function $R_{k}(o,\phi_{k})$ scores each output,where $\phi_{k}$ denotes task-specific information required for the evaluation, such as reference answers, test cases, or comparison outputs required for evaluation. The policy $\pi_{\theta}$ is updated to maximize expected return while controlling divergence from a reference policy.
Let
$\mu_{k}=\text{mean}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},\quad\sigma_{k}=\text{std}\big{(}\{R_{k}(o_{k,i},\phi_{k})\}_{i=1}^{G}\big{)},$ (3)
we define the policy ratio $\rho_{k,i,t}$ and the normalized advantage estimate $\hat{A}_{k,i,t}$ as
$\rho_{k,i,t}=\frac{\pi_{\theta}(o_{k,i,t}\mid q_{k},o_{k,i,<t})}{\pi_{\theta_{\text{old}}}(o_{k,i,t}\mid q_{k},o_{k,i,<t})},$ $\quad\hat{A}_{k,i,t}=\frac{R_{k}(o_{k,i},\phi_{k})-\mu_{k}}{\sigma_{k}}.$ (4)
This allows us to write the MT-GRPO objective as
$\begin{split}&\mathcal{J}_{\text{MT-GRPO}}(\theta)=\mathbb{E}_{k\sim K,q_{k}\sim\mathcal{D}_{k},\{o_{k,i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot\mid q_{k})}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{k,i}|}\sum_{t=1}^{|o_{k,i}|}\left\{\min\left[\rho_{k,i,t}\hat{A}_{k,i},\text{clip}\left(\rho_{k,i,t},1{-}\epsilon,1{+}\epsilon\right)\hat{A}_{k,i}\right]-\beta_{k}\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (5)
where
$\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]=\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-\log\frac{\pi_{ref}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}-1.$ (6)
4.3 Task Scheduling
We consider two strategies for optimizing a shared policy across multiple tasks: curriculum learning and joint training. Both utilize task-specific rewards but differ in scheduling. To guide task ordering, we adapt a Backward Transfer (BWT) metric to quantify forgetting in our multi-task setting. Following Lopez-Paz & Ranzato (2017), we define $BWT_{j}=P_{after,j}-P_{base,j}$, where $P_{base,j}$ is the performance on task $j$ from the original base model and $P_{after,j}$ is the performance on task $j$ after training on subsequent tasks. A negative $BWT_{j}$ indicates that performance on task $j$ degrades due to learning new tasks, measuring the extent of forgetting.
Curriculum Learning:
Prior work in curriculum learning prioritizes task complexity (e.g., easy to hard) for more effective learning (Bengio et al., 2009; Parashar et al., 2025). We instead adopt a forgetting-based curriculum that orders tasks according to their average forgettability in multi-task learning. Tasks with the least negative or even positive backward transfer are introduced earlier, serving as stable foundations for subsequent tasks that are more prone to catastrophic forgetting. In our setup, this coincides with increasing supervision subjectivity. This forgetting-based ordering reduces overall forgetting and improves generalization by minimizing destructive interference across tasks.
Joint Training:
All tasks are sampled concurrently from a shared distribution. While this promotes exposure, it introduces three key challenges: (1) simpler tasks may dominate without adaptive sampling due to the tendency of reward hacking in GRPO; (2) reward delays vary, requiring asynchronous updates; and (3) Kullback–Leibler (KL) regularization demands tuning per task due to varying sensitivities.
In both regimes, we apply MT-GRPO with normalized advantages and task-specific KL coefficients. At each step, a task $T_{k}$ is selected, completions are generated by the current policy $\pi_{\theta}$, and updates are applied based on the corresponding reward signal.

### v3

# Omni-Thinker: Scaling Multi-Task RL in LLMs with Hybrid Reward and Task Scheduling

@@ Abstract
The pursuit of general-purpose artificial intelligence depends on large language models (LLMs) that can handle both structured reasoning and open-ended generation. We present Omni-Thinker, a unified reinforcement learning (RL) framework that scales LLMs across diverse tasks by combining hybrid rewards with backward-transfer–guided scheduling. Hybrid rewards integrate rule-based verifiable signals with preference-based evaluations from an LLM-as-a-Judge, enabling learning in both deterministic and subjective domains. Our scheduler orders tasks according to accuracy backward transfer (BWT), reducing forgetting and improving multi-task performance. Experiments across four domains show gains of $6.2\%$ over joint training and $12.4\%$ over model merging. Moreover, we demonstrate that simple assumptions on accuracy transfer yield accurate predictions of curriculum outcomes, with entropy dynamics explaining deviations due to generative tasks. These findings underscore the importance of BWT-aware scheduling and hybrid supervision for scaling RL-based post-training toward general-purpose LLMs.
@@ 1 Introduction
Figure 1: Performance across four task domains, comparing Joint Training and Curriculum Learning against baselines including SFT and Model Merging. Curriculum Learning achieves the strongest results, showing that controlling how tasks are scheduled is crucial for effective multi-task learning.
@@
Reinforcement learning (RL) has become an effective approach for improving large language models (LLMs) (Hurst et al., 2024; Liu et al., 2024; Dubey et al., 2024; Yang et al., 2024), particularly in structured domains such as math and coding where verifiable, rule-based rewards are available (Guo et al., 2025; Luo et al., 2025; Kimi-Team et al., 2025). Methods such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) show that even coarse learning signals can steer LLMs toward structured, chain-of-thought reasoning. However, most RL methods remain tailored to deterministically verifiable tasks, limiting their utility in open-ended domains such as general QA and creative writing. Moreover, training LLMs across multiple tasks remains challenging because it requires optimizing for diverse forms of feedback signals, including binary correctness checks in structured tasks and subjective, preference-based judgments in generative ones.
@@
We address this challenge with Omni-Thinker, a unified RL framework that enables LLMs to learn from both rule-based and generative supervision under a single policy. Building on Reinforcement Learning with Verified Reward (RLVR), our method integrates symbolic verifiers with LLM-as-a-Judge evaluations (Zheng et al., 2023; Zhang et al., 2025) to handle subjective tasks. Our curriculum is forgetting-aware; it is guided by backward transfer (BWT), where BWT denotes test-performance backward transfer computed on a normalized, task-specific test metric. Ordering task training according to this signal yields effective curricula across heterogeneous domains. We show that the final accuracy of model after curriculum learning is well predicted by forgettability ranking, even under simplifying assumptions. Empirically, we observe complementary entropy dynamics, fine-tuning on creative writing tends to increase the model’s output entropy, whereas training on verifier-supervised, structured tasks tends to decrease it; this trend is consistent with our BWT-guided choice to train structured tasks before open-ended ones. Across four domains, Omni-Thinker improves generalization while reducing forgetting, with average gains of 6.2% over joint multi-task training and 12.4% over model merging, respectively.
Our key contributions are threefold. (1) We present Omni-Thinker, a unified framework that trains a single policy across four diverse domains, using hybrid verifiable and preference-based rewards. (2) We develop a forgetting-aware curriculum based on backward transfer (BWT) linear ordering maximization over task-specific test performance to reduce forgetting, outperforming joint multi-task training and model merging. (3) We empirically analyze training dynamics through the lens of entropy, revealing that structured domains (math, coding) systematically decrease output entropy while open-ended domains (creative writing) increase it, thereby providing an explanatory link between entropy evolution and the effectiveness of BWT-guided curricula.
2 Framework Overview
We introduce Omni-Thinker as a unified reinforcement learning framework for large language models that integrates hybrid rewards with task scheduling guided by backward transfer. Unlike prior approaches that separate reasoning and generative domains, Omni-Thinker maintains a single policy across heterogeneous tasks, including Math, Coding, General QA, and Creative Writing, while dynamically ordering training to minimize forgetting. The framework is instantiated using Multi-Task GRPO, augmented with both symbolic verifiers and LLM-as-a-Judge supervision, and a curriculum determined by accuracy- and entropy-based backward transfers.
2.1 Notation and Training Objective
We give ourselves a vocabulary $\mathcal{V}$ with a special end-of-sequence token $\mathrm{eos}$. The set of finite sequences of tokens is denoted $\mathcal{V}^{*}$; for any sequence $o\in\mathcal{V}^{*}$, its length is denoted $|o|$ and we say that $o$ is complete if $o_{|o|}=\mathrm{eos}$. A model, parameterized by $\theta$, defines a conditional distribution $\pi_{\theta}(y_{t}\mid y_{<t})$ for any given sequence of tokens $(y_{t})_{t\in\mathbb{N}}$. It induces a policy $\pi_{\theta}^{\otimes}$ on token sequences defined by $\pi_{\theta}^{\otimes}(o\mid q,o_{<t_{0}}):=\prod_{t=t_{0}}^{|o|}\pi_{\theta}(o_{t}\mid q,o_{<t})$. We adopt a multi-task RL (MTRL) formulation: a task is a couple $T=(\mathcal{D},R)$ where $\mathcal{D}$ is a dataset of prompts and $R(q,o)$ is a task-specific reward function. Given a set of $K$ tasks $\mathcal{T}=\{T_{1},\ldots,T_{K}\}$, the goal is to learn a unified policy $\pi_{\theta}$ that maximizes the expected reward over the task distribution:
$\max_{\theta}\;\mathcal{J}(\theta)=\mathbb{E}_{(\mathcal{D},R)\sim P(\mathcal{T})}\left[\mathbb{E}_{q\sim\mathcal{D},o\sim\pi_{\theta}^{\otimes}(\cdot\mid q)}\left[R(o)\right]\right],$ (1)
where $P(\mathcal{T})$ is a task sampling distribution, which determines task exposure during training.
In order to train $\pi_{\theta}$ to maximize the objective $\mathcal{J}$, we extend the GRPO (Guo et al., 2025) algorithm to the multi-task setting by jointly optimizing over task-specific reward signals and reference policies. For each input prompt $q$, GRPO samples a group of outputs $\{o_{q,1},o_{q,2},\cdots,o_{q,G}\}$ from the old policy $\pi_{\theta_{old}}$. A task-specific reward function $R_{k}(q,o)$ scores each output. The policy $\pi_{\theta}$ is updated to maximize expected return while controlling divergence from a reference policy.
We define the policy ratio $\rho_{q,i,t}$ and the normalized advantage estimate $\hat{A}_{q,i,t}$ as follows:
$\mu_{q}=\text{mean}\big(\{R_{k}(q,o_{q,i})\}_{i=1}^{G}\big),$ $\sigma_{q}=\text{std}\big(\{R_{k}(q,o_{q,i})\}_{i=1}^{G}\big),$ (2)
$\rho_{q,i,t}=\frac{\pi_{\theta}(o_{q,i,t}\mid q,o_{q,i,<t})}{\pi_{\theta_{\text{old}}}(o_{q,i,t}\mid q,o_{q,i,<t})},$ $\hat{A}_{q,i,t}=\frac{R_{k}(q,o_{q,i})-\mu_{q}}{\sigma_{q}}.$ (3)
This allows us to write the MT-GRPO objective as
@@ 2.1 Notation and Training Objective
$\begin{split}&\mathcal{J}_{\text{MT-GRPO}}(\theta)=\mathbb{E}_{k\sim K,q\sim\mathcal{D}_{k},\{o_{q,i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}^{\otimes}(\cdot\mid q_{k})}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{q,i}|}\sum_{t=1}^{|o_{q,i}|}\left\{\min\left[\rho_{q,i,t}\hat{A}_{q,i},\text{clip}\left(\rho_{q,i,t},1{-}\epsilon,1{+}\epsilon\right)\hat{A}_{q,i}\right]-\beta_{k}\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]\right\},\end{split}$ (4)
where
$\mathbb{D}_{KL}\left[\pi_{\theta}||\pi_{ref}\right]=\frac{\pi_{ref}(o_{q,i,t}|q,o_{q,i,<t})}{\pi_{\theta}(o_{q,i,t}|q,o_{q,i,<t})}-\log\frac{\pi_{ref}(o_{q,i,t}|q,o_{q,i,<t})}{\pi_{\theta}(o_{q,i,t}|q,o_{q,i,<t})}-1.$ (5)
2.2 Hybrid Rewards
We design a hybrid reward system that unifies reinforcement learning across both structured reasoning tasks and open-ended generative domains.
@@ 2.2 Hybrid Rewards
Together, these components form a unified hybrid reward scheme: verifiable rewards ensure correctness where possible and generative-based signals cover subjective domains. This design enables reinforcement learning to scale across diverse tasks, from reasoning to open-ended generation.
2.3 Joint training and Curriculum Learning
In practice, a maximization step of the training objective $\mathcal{J}_{\text{MT-GRPO}}$ requires a batch $B$ of prompts sampled from $\bigcup_{k=1}^{K}\mathcal{D}_{k}$ then sampling a batch of outputs $\{o_{q,i}\}_{i=1}^{G}$ for each $q\in B$. A multi-task schedule is defined as a sequence of batches $(B_{s})_{s=1}^{s_{\max}}$ such that $\forall s\neq s^{\prime},B_{s}\cap B_{s^{\prime}}=\emptyset$ and $\bigcup_{s=1}^{s_{\mathrm{max}}}B_{s}=\bigcup_{k=1}^{K}\mathcal{D}_{k}$.
Two special cases are considered: Joint Training and Curriculum Learning. Joint Training consists in sampling each batch $B$ uniformly at random among all samples (without replacement), disregarding their corresponding tasks: $\forall s,B_{s}\sim\mathcal{U}\left(\bigcup_{k=1}^{K}\mathcal{D}_{k}\setminus\bigcup_{s^{\prime}<s}B_{s^{\prime}}\right)$. Curriculum Learning on the other hand consists of pure batches chosen from the same task until exhaustion of the task dataset. By pure, we mean that each batch is derived from only one task dataset: $\forall s,B_{s}\sim\mathcal{U}\left(\mathcal{D}_{k_{s}}\setminus\bigcup_{s^{\prime}<s}B_{s^{\prime}}\right)$ for some task schedule $(k_{s})_{s\in\{1,\cdots,s_{\mathrm{max}}\}}$. A Curriculum is described by a permutation $\sigma\in\mathfrak{S}_{K}$ of the tasks, with $\mathfrak{S}_{K}$ the set of permutation of $\{1,\cdots,K\}$.
3 Backward Transfer for Task-scheduling
We intend to use Backward Transfers (BWT) to guide our choice of curriculum. Following Lopez-Paz & Ranzato (2017) it is defined as follows.
Definition 1 (Backward Transfer Matrix).
Let $\theta_{0}$ be a set of initial parameters of a model $\pi_{\theta}$ and let $\theta(\theta_{0},T)$ be the set of parameters obtained after training $\pi_{\theta}$ on task $T$ starting from $\theta_{0}$. Write $\mathrm{Acc}(\theta,T)$ the accuracy of model $\pi_{\theta}$ on task $T$. The backward transfer matrix is defined by
$\mathrm{BWT}_{ij}(\theta_{0}):=\log\mathrm{Acc}({\theta(\theta_{0},T_{j})},T_{i})-\log\mathrm{Acc}({\theta_{0}},T_{i}).$ (6)
Input: $\mathrm{BWT}\in\mathbb{R}^{K\times K}$. $a_{\text{init}}:=\left(\mathrm{Acc}(\theta_{0},T_{k})\right)_{k=1}^{K}\in\mathbb{R}^{K}$. Curriculum $\sigma\in\mathfrak{S}_{K}$
$a\leftarrow a_{\text{init}}$;
for $j=1$ to $K$ do
 $k\leftarrow\sigma(j)$;
 $a_{k}\leftarrow a_{\text{init},k}$;
 for $i=1$ to $K$ do
    $a_{i}\leftarrow a_{i}\times\exp(\mathrm{BWT}_{ik})$;
   end for
end for
return $a$;
Algorithm 1 Final Accuracy under Assumptions 1 and 2
3.1 A priori prediction of terminal accuracies under constant BWT
Our goal is to choose a curriculum order $\sigma$ a priori by predicting the terminal per–task accuracies without training all permutations. We propose a simple predictive model in which (i) inter–task backward transfers are treated as constant in log–accuracy, and (ii) training on the full dataset of a task saturates its self–accuracy. Under these assumptions, terminal accuracies for any order $\sigma$ become computable from quantities measured once at initialization.
Setup.
Using notations from Section 2.3, let $\theta_{0}$ denote the parameters of the pre–trained model $\pi_{\theta}$, and let $\theta_{s}$ be the parameters after $s$ optimization steps following a curriculum order $\sigma\in\mathfrak{S}_{K}$.
Assumption 1 (Constant off–diagonal BWT in log–accuracy).
For all $i\neq j$ and all optimization steps $s$ along the schedule,
$\mathrm{BWT}_{ij}\!\left(\theta_{s}\right)=\mathrm{BWT}_{ij}\!\left(\theta_{0}\right).$ (7)
Assumption 2 (Task–wise saturation).
Let $\{B_{s}\}$ be the sequence of mini–batches processed along the schedule. If, between steps $s_{1}$ and $s_{2}$, the full dataset $\mathcal{D}_{k}$ of task $T_{k}$ has been seen, then accuracy saturates on task $T_{k}$ to the same accuracy as training from $\theta_{0}$:
$\bigcup_{s=s_{1}}^{s_{2}}B_{s}=\mathcal{D}_{k}\quad\Rightarrow\quad\mathrm{Acc}\!\left(\theta_{s_{2}},T_{k}\right)=\mathrm{Acc}\!\big(\theta(\theta_{0},T_{k}),\,T_{k}\big).$ (8)
Theorem 1.
Under Assumptions 1 and 2, for any curriculum order $\sigma\in\mathfrak{S}_{K}$ starting from $\theta_{0}$, the terminal accuracies $\{\mathrm{Acc}(\theta_{s_{\max}},T_{k})\}_{k=1}^{K}$ given the initialization accuracies $\{\mathrm{Acc}(\theta_{0},T_{k})\}_{k=1}^{K}$, the $\mathrm{BWT}(\theta_{0})$ and a curriculum $\sigma$ is exactly the output of Algorithm 1
Reasonableness and limitations.
Assumption 1 abstracts away known drivers of transfer, such as domain overlap, stochasticity, and entropy evolution, thus is a toy yet useful approximation for a priori curriculum selection. See Sections 5.2 and 5.3 for a discussion of a correction coming from entropy. Working in log–accuracy space keeps accuracies positive but does not eliminate the risk of unrealistic growth over long curricula (e.g., predictions exceeding $1$ in accuracy). Assumption 2 is reasonable when each task dataset is sufficiently large and optimization keeps the model in a well–conditioned regime, conditions we satisfy in our experiments. Together, these assumptions yield a tractable predictor that captures coarse curriculum effects while remaining simple enough to evaluate without exhaustive training.
3.2 Curriculum Choice via Linear Ordering Maximization
Algorithm 1 admits the following closed form for the predicted terminal accuracies:
$\log\mathrm{Acc}(\sigma)-\log\mathrm{Acc}(\mathrm{Id})=\sum_{i<j}\bigl(\Phi_{\sigma}^{-1}\,\mathrm{BWT}\,\Phi_{\sigma}\bigr)_{ij},\quad\text{with}\quad\Phi_{\sigma,ij}=\mathbf{1}_{i=\sigma(j)}.$ (9)
Input: BWT matrix $\mathrm{BWT}\in\mathbb{R}^{K\times K}$
$\sigma\leftarrow$ empty list;
while there are unvisited tasks do
 $k^{*}\leftarrow\arg\max_{k\notin\sigma}\ \sum_{i\notin\sigma\cup\{k\}}\mathrm{BWT}_{ik}$;
   append $k^{*}$ to $\sigma$;
end while
return $\sigma$;
Algorithm 2 Greedy BWT-LOM Curriculum
In words, curriculum reordering acts by permuting the BWT matrix with $\Phi_{\sigma}$, and the gain relative to the identity schedule is simply the sum of the upper–triangular entries of the permuted matrix.
Given an aggregated score of the form
$\mathcal{S}:=\sum_{T\in\mathcal{T}}\alpha_{T}\log\mathrm{Acc}(\theta,T),$ (10)
identifying the best task order amounts to solving a Linear Ordering Problem (LOP). see (Floudas & Pardalos, 2008) for an overview. This problem is known to be NP–hard, but for small numbers of tasks ($K$) it can be solved exactly. For larger $K$, a wide range of approximation algorithms and heuristics exist. A simple heuristic is to rank tasks by a forgettability score: $F_{k}:=\alpha_{k}\sum_{i\neq k}\mathrm{BWT}_{ik}$. Intuitively, ordering tasks by decreasing $F_{k}$ prioritizes those that exert the least destructive interference on others (or even provide positive transfer), thereby reducing overall forgetting. Our curriculum orders tasks by decreasing column mean of BWT.